OpenStudy (anonymous):

(2x -5)/(x-2) -2 = 3/(x+2) I have been trying to figure this out for the past half hour I am losing it here =(.

OpenStudy (anonymous):

\[\frac{ 2x+5 }{ x-2 }-2=\frac{ 3 }{ x+2}\]

OpenStudy (anonymous):

is that how it looks?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your trying to find x?

OpenStudy (anonymous):

yes. And they want to show work, problem is every time I try to solve I come up with some thing different.

OpenStudy (anonymous):

meant 2x-5 up there not 2x+5

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

I know that because I used an algebraic calculator but I can't find a way to solve the equation to show that.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

he's the best at explaining stuff

jimthompson5910 (jim_thompson5910):

The LCD is (x-2)(x+2), so if you multiply everything by this LCD, you'll clear out the fractions (2x -5)/(x-2) -2 = 3/(x+2) (x-2)(x+2)*(2x -5)/(x-2) -2(x-2)(x+2) = (x-2)(x+2)*3/(x+2) (x+2)(2x-5) - 2(x-2)(x+2) = 3(x-2) 2x^2 - 5x + 4x - 10 - 2(x^2 - 4) = 3(x-2) 2x^2 - 5x + 4x - 10 - 2x^2 + 8 = 3x - 6 -x - 2 = 3x - 6 I'll let you finish up

jimthompson5910 (jim_thompson5910):

Make sure you check any possible answers you get

OpenStudy (anonymous):

wow! Thank you so much! That helps a lot

OpenStudy (anonymous):

ok what about this one? 4/(x+3) -1/2 = 3/(2x+6) I don't know how to start without an LCD

jimthompson5910 (jim_thompson5910):

2x+6 factors to 2(x+3)

jimthompson5910 (jim_thompson5910):

so the LCD is 2(x+3)

OpenStudy (anonymous):

correction 2/(x+3) -1/2 = 3/2(x+3)

OpenStudy (anonymous):

so 2/x+3 * 2(x+3) cancels out? Sorry I typed it wrong at first.

jimthompson5910 (jim_thompson5910):

yes multiplying 2/x+3 * 2(x+3) out will give you 2*2 = 4

jimthompson5910 (jim_thompson5910):

do the same for every other term

OpenStudy (anonymous):

x = -6?

jimthompson5910 (jim_thompson5910):

did you get just that one answer? if so, then it's incorrect

OpenStudy (anonymous):

-1/2* 2(x +3) -1 (x + 3) -x -3 = 3 -x = 6 x = -6

jimthompson5910 (jim_thompson5910):

you forgot the 4 we found earlier

jimthompson5910 (jim_thompson5910):

it should be 4 - 1(x+3) = 3

jimthompson5910 (jim_thompson5910):

solve that and you'll get your answer

OpenStudy (anonymous):

*headdesk* wow. Thanks! I've been awake too long...

jimthompson5910 (jim_thompson5910):

that's ok, don't take it out on your head or the desk though lol you should try to get some sleep if you've been up for a while

OpenStudy (anonymous):

finals week gotta get it all done =(

jimthompson5910 (jim_thompson5910):

i gotcha, well i wish you the best of luck